OpenStudy (anonymous):

graph the equation by plotting three points -5x-3y=-7

OpenStudy (anonymous):

can someone please help me

OpenStudy (phi):

can you change this to y = mx +b format ? the first step is add +5x to both sides. what do you get ?

OpenStudy (anonymous):

10x+2y=-2

OpenStudy (anonymous):

phi

OpenStudy (phi):

remember what we are trying to do: get y by itself. that should be a clue. -5x-3y=-7 now add +5x to both sides +5x - 5x - 3y = 5x -7 on the left side you have 5 x's take away 5 x's. how many x's do you have left ?

OpenStudy (anonymous):

0x

OpenStudy (phi):

yes, so +5x - 5x - 3y = 5x -7 simplifies to -3y = 5x-7 already this equation looks familiar... but to continue divide both sides (and *all* terms) by -3. what do you get ? (and you already did this problem, so you should try to get it right)

OpenStudy (anonymous):

1y=6x-7

OpenStudy (phi):

to divide by -3, just write /-3 next to each term in -3y = 5x-7 can you do that ?

OpenStudy (anonymous):

-3y/-3=5x/-3-7/-3

OpenStudy (phi):

yes, now simplify -3y/-3 is 1y (because -3/-3 is 1) or just y 5x/-3 is also written as (5/-3) x and -7/-3 is 7/3 you get y = (5/-3)x + 7/3 or y= (-5/3) x + 7/3 both are the same thing. or using the equation editor \[ y = \frac{-5}{3} x + \frac{7}{3} \]

OpenStudy (phi):

didn't you already do this problem ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

its it going to be the same points

OpenStudy (phi):

you could find 3 different points... lines have an infinite number of points... on the other hand, it is easier to use the first 3 you found

OpenStudy (anonymous):

this one (-1,4) (2,-1) (5,-6)

OpenStudy (phi):

yes

OpenStudy (anonymous):

thank you could you help me with another one

OpenStudy (phi):

did you post it ?

OpenStudy (anonymous):

im doing it now